OpenStudy (anonymous):

Show that the sum of the x-,y-&z- intercepts of any tangent plane to the surface √x+ √(y )+ √z= √c is a constant.

OpenStudy (anonymous):

tangent plane equation: \[\frac{\partial f}{\partial x}x+\frac{\partial f}{\partial y}y+\frac{\partial f}{\partial z}z=0\] where f(x,y,z) is surface equation Later find it's intercepts by setting x=0,y=0 for z-inrcp. And similar for the rest

OpenStudy (anonymous):

is z intercept is 0.5?

OpenStudy (anonymous):

dind't calculate it.

OpenStudy (anonymous):

am i right?